Tariyah:

Select two ratios that are equivalent to 3:1, (Choice A) A 12:48 (Choice B) B 6:18 (Choice C) C 4:12 (Choice D) D 3:1 (Choice E) E 1:4

Isolation:

3:1 This is like saying \[\frac{ 3 }{ 1 }\] Let's look at the first one 12:48, like \[\frac{ 12 }{ 48 }\] Do you think that \[\frac{ 3 }{ 1 } = \frac{ 12 }{ 48 }\] ?

Solaimon15:

u cant give direct answers my g

Tariyah:

its not a,d

Solaimon15:

u sold

Isolation:

Are you sure you wrote the question correctly? because looking at it like this... only one is correct

Tariyah:

i did thats y im confused

Isolation:

Can you screenshot the question maybe?

Tariyah:

Solaimon15:

srry tariyah my g sold

Tariyah:

I chose A and E and I got to correct

Solaimon15:

good

Tariyah:

thx for the help

Isolation:

Oh, you meant 3:12

Tariyah:

yelp

Solaimon15:

ight

Solaimon15:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Tariyah yelp \(\color{#0cbb34}{\text{End of Quote}}\) take care